OpenStudy (anonymous):

Graph the following expression on the number line by placing the dot in the proper location. |2 - x | = 4

OpenStudy (anonymous):

For this to be true 2 - x = 4 OR 2 - x = -4 x = -2 x = 6 Put dots on -2 and 6 on your number line.

OpenStudy (jdoe0001):

\(\large |2 - x | = 4\implies \begin{cases} +(2 - x ) = 4\implies 2 - x = 4\\ \quad \\ -(2 - x ) = 4\implies 2 - x =-4 \end{cases}\)